OpenStudy (anonymous):

Plz help Find the length of the arc of the parabola x^2=8y from x=0 and x=4 Waiting ASAP

OpenStudy (sirm3d):

use this formula for arclength\[\Large L=\int_{x_0}^{x_1}\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]